A man walks up a stationary escalator in 90sec.when this man stands on moving escalator he goes up in 60sec.the time taken by the man to walk up the moving escalator is

Grade:11

1 Answers

Vikas TU
7 years ago
walking up when stationary = 90 sec.
l = v1*90
when escalator is moving,
l = v2*60
when botth are moving,
l = (v1 + v2)*t
solving the 3 eqns.
v2 = 1.5v1
and
90v1 = (2.5v1)t
t = 90/2.5 = > 900/25 = > 180/5 = > 36 s.